OpenStudy (anonymous):

(a + b - c)(a + b + c) as an answer i got a^2 +b^3 - c^2 is that right?

OpenStudy (anonymous):

You have: \[a(a+b+c)+b(a+b+c)-c(a+b+c)=a^2+ab+ac+ab+b^2+bc-ca-cb-c^2\] \[=a^2+2ab+b^2-c^2\]